### v1

# Gas Cost Analysis of Proxy and Diamond Patterns: Towards Trusted Smart Contract Engineering in EVM Blockchains

Abstract
Blockchain applications are witnessing rapid evolution, necessitating the integration of upgradeable smart contracts. Software patterns have been proposed to summarize upgradeable smart contract best practices. However, research is missing on the comparison of these upgradeable smart contract patterns, especially regarding gas costs related to deployment and execution. This study aims to provide an in-depth analysis of gas costs associated with two prevalent upgradeable smart contract patterns: the Proxy and diamond patterns. The Proxy pattern utilizes a Proxy pointing to a logic contract, while the diamond pattern enables a Proxy to point to multiple logic contracts. We conduct a comparative analysis of gas costs for both patterns in contrast to a traditional non-upgradeable smart contract. We derive from this analysis a theoretical contribution in the form of two consolidated blockchain patterns and a corresponding decision model. By so doing we hope to contribute to the broader understanding of upgradeable smart contract patterns.
Keywords:
Blockchain Software Patterns Upgradeable Smart Contracts Proxy Pattern Diamond Pattern
1 Introduction
In the world of blockchain technology, smart contracts stand as keystones to orchestrate reliably and securely digital transactions [20]. Smart contracts are particularly important on Ethereum and similar blockchain platforms, where they are used in diverse areas including digital finance, identity management, or industrial traceability.
As blockchain applications evolve, the imperative for smart contracts to be not only secure but also upgradeable becomes apparent [10]. In this context, a set of upgradeable smart contract patterns, including the Proxy, and Diamond, have surfaced to answer the lack of adaptability of classical smart contracts whose logic cannot be changed once deployed [23, 27]. The Proxy pattern, wielding a proxy contract to delegate calls to an implementation contract, offers a flexible and upgradeable alternative. The Diamond pattern, as proposed in EIP-2535, addresses concerns like maximum contract size, ensuring enhanced maintainability and versioning through multiple implementation contracts.
Research has focused on identifying upgradeable smart contracts pattern families, pointing towards Proxy and Diamond-based strategies [22, 16, 23, 27]. However, the papers do not provide an in-depth analysis of the functional and non-functional properties of these patterns, nor gas costs behaviors. Hence, a research gap is identified regarding a thorough study of the Proxy and Diamond patterns especially in terms of a gas costs analysis.
To address this issue, this paper aims to answer the following research questions: (RQ1) How do Proxy and Diamond patterns differ in terms of gas consumption, scalability, and ease of use? and (RQ2) What implications do these differences hold for developing blockchain applications?
In this paper, we contribute to the literature through a unified approach to compare gas costs in upgradeable smart contracts. We leverage this methodology to provide a comparative gas cost analysis of the Proxy and Diamond patterns in EVM blockchains, compared against a monolithic non-upgradeable smart contract, which is used as baseline. Based on these results, we derive a theoretical contribution in the form of two smart contract patterns adhering to the Alexandrian form format following the standard proposed by Christopher Alexander [2]. These patterns include the results of our comparative analysis and contribute to the broader understanding of upgradeable smart contract patterns and their use in blockchain applications. Based on these patterns, we also propose a decision model for using these patterns that highlights functional and non-functional properties involved in each decision design path.
The remainder of this paper is structured as follows. Section 2 introduces key concepts related to smart contracts and blockchain patterns, and section 3 presents studies already made on linked concepts. Section 4 presents the method used to carry our comparative analysis on Proxy versus Diamond gas costs. Section 5 presents the results of the tests made on the different patterns. Section 6 leverages these findings to propose two consolidated Proxy and Diamond patterns as well as a decision model for using these patterns. Section 7 finally concludes the paper with a summary and a discussion of the results and some considerations for future work.
2 Background
This section introduces key concepts that will be used in the sequel of this paper revolving around smart contracts and related concepts such as upgradability patterns, EVM, gas and storage.
2.1 Smart Contracts
A smart contract is a program hosted on a blockchain network [30]. When a smart contract executes, its updated state (or storage) is registered into a transaction that will be stored in the blockchain ledger making it immutable and tamper-proof [5]. More precisely, the final validated results are stored in a Merkle Patricia trie whose nodes correspond to an account or a smart contract. A smart contract comprises both variables and functions. Smart contract functions can execute arbitrary code, access the state of the variables and optionally update them. The default function’s mutability gives right to read and modify state variables. However, function mutability can be constrained to add more security when it comes to accessing these variables. On the one hand, a pure function cannot read nor modify the state of the contract; it is only used to compute a value, often using the parameters passed to it. On the other hand, a view function can read the state of the contract, but cannot modify it.
Smart contracts are immutable, meaning that once deployed, they cannot be modified [31, 9]. This is a security feature, as it prevents malicious actors from modifying the code, so that the contract can be trustable and unbreakable. However, this also means that if a bug is found in the code, it cannot be fixed, and the only solution is to deploy a new contract. Also, if a new feature is added to the contract, the only way to do it is to deploy a new contract. This is a problem for users, as they have to migrate to the new contract, and for developers, as they have to maintain multiple contracts. To address this issue, several patterns have emerged [10, 23], which are discussed later in the paper.
2.2 Gas and storage in the Ethereum Virtual Machine (EVM)
The EVM is a Turing-complete virtual machine that runs on every node of the Ethereum network and other EVM-based blockchains. It provides a secure and isolated environment for smart contracts execution [18]. Gas is the unit of computation in the EVM, and every operation performed by the EVM has a gas cost associated with it [17]. The user pays this cost in the form of gas fees, to the miner/validator who executes the smart contract. It is the miner’s incentive to execute the smart contract and record the results on the blockchain, but also a way to prevent spam and denial of service attacks.
Storage in the EVM is linked to the concept of gas, as every operation involving storage – whether it is writing new data or modifying existing data – incurs a gas cost. This cost is proportional to the storage resources consumed, reflecting the principle that the more network resources (like memory and storage space) a transaction uses, the more it needs to compensate the network. Every smart contract has its own storage space, which is isolated and theoretically unlimited.
The EVM utilizes a key-value store for data storage, with these key-value pairs referred to as storage slots. A slot’s key is the hash of the contract’s address and the hash of its value. A value is a 32-byte (or 256-bit) word, also called an item. Data smaller than 32 bytes can be packed into a single slot, but if it is larger, the transaction will be reverted [14]. Some types of data are stored in multiple slots, such as arrays and mappings which are stored in multiple contiguous slots [15]. Strings are also stored in multiple slots, where one slot stores the length of the string, and the other slots store each 32-byte chunk of the string [13]. A struct, which is a collection of variables, is stored in a single slot if it fits, otherwise, it is stored in multiple contiguous slots. The gas cost of writing to storage increases with the number of slots written to. This means that writing multiple variables, if they fit within a single slot, incurs the same gas cost as writing a single variable.
3 Related Work
Software patterns offer a promising approach to this challenge, providing well-tested solutions for frequently encountered application development use cases [3]. A family of software patterns focus on blockchain patterns, which include data management, domain-based, or smart-contract related best practices [28]. In the subcategory of smart contract patterns stand upgradability patterns, which focus on a set of recurring software design approaches to maintain decentralized applications when adding new features or fixing issues are needed [10]. Two recurring patterns emerge through the literature, the Proxy and Diamond patterns [22, 16, 23, 27]
The Proxy pattern enables smart contract updates without changing the contract’s address or requiring data migration[26]. Comprising two contracts – the user-interacted proxy and the logic-holding implementation – the proxy forwards calls to the implementation via delegated calls. Shared storage ensures that if the implementation is updated, the storage persists. To update the implementation, a new contract is deployed, and the proxy is directed to it, eliminating the need for user migration and preventing data loss. Compound, a key DeFi protocol, exemplifies this pattern through multiple upgrades, including Compound III  [^1] https://compound.finance/. Similarly, USDC, a widely-used stable coin, leverages the Proxy pattern for notable upgrades like the transition to USDC 2.0, showcasing its adaptability in the evolving digital currency landscape  [^2] https://www.coinbase.com/blog/usdc-v2-upgrading-a-multi-billion-dollar-erc-20-token.
Figure 1: Diamond Pattern Diagram
The Diamond pattern, depicted in Figure 1 is a more upgradeable version of the Proxy pattern. This pattern solves the maximum contract size problem which is a limitation of the Proxy pattern. Indeed, logic can be separated into small contracts referred to as facets. A main contract, referred to as the implementation or diamond contract, points at the different facets to retrieve the applicable logic. The diamond contract can be upgraded by adding, replacing or removing facets. The Diamond pattern is being utilized by a diverse array of projects, as documented in Nick Mudge’s Awesome Diamonds repository[25]. For example, Aavegotchi, a Non-Fungible Token (NFT) based gaming protocol, employs a single Diamond pattern with eight distinct facets  [^3] https://www.aavegotchi.com/. Each facet serves a specific purpose, contributing to the protocol’s overall functionality. This implementation illustrates how the Diamond pattern can efficiently organize and manage complex functionalities within a single contract structure, an essential feature for sophisticated blockchain applications like gaming protocols.
Two studies focus on gas cost efficiency strategies in smart contracts development. Zarir et al.’s work [33] focuses on transaction parameters rather than architectural design, while the study by Di et al. [12] identifies smart contract coding metrics impacting gas costs but do not specifically focus on upgradeable smart contracts, which possess unique functionalities like proxy pointers and proxy contract management.
A set of studies provide insight on upgradeable smart contract patterns. Kannengiesser et al. [20] conduct a study on key smart contract development challenges across various distributed ledger technology (DLT) protocols, including Ethereum, Hyperledger, and EOSIO. They highlight upgradeability as a significant challenge and reference two upgradeable smart contract patterns, namely the Diamond (referred to as the Façade pattern) pattern and the Proxy pattern. However, the paper lacks an in-depth analysis of these patterns. Two papers [22, 16] identify the Proxy pattern but do not detail the forces, advantages, drawbacks, or gas costs considerations of this pattern. There is no mention of the Diamond pattern. Two other works present a set of upgradeable smart contract patterns, including the Proxy and Diamond patterns [23, 27]. However, the papers do not provide an in-depth analysis of the functional and non-functional properties of these patterns, nor gas costs behaviors.
In summary, a research gap is identified regarding a thorough study of the Proxy and Diamond patterns especially in terms of a gas costs analysis.
4 Methodology
The methodology section of this paper details the approach employed for comparing the Proxy and Diamond patterns. This involves the definition of a baseline scenario and the development of a gas consumption evaluation test bench.
4.1 Protocol
The initial smart contract deployment can be one of the most expensive in terms of gas costs in a smart contract life cycle. Therefore, assessing the deployment costs associated with various patterns is essential. Additionally, upgrades often necessitate deploying additional smart contracts, making it crucial to compare the deployment costs of each subsequent upgrade.
To do so, we propose to develop a file notarization scenario which is a standard use of smart contracts to ensure the integrity of critical data such as diploma or scientific workflows [6, 11, 21]. The scenario will be implemented using the Proxy, and Diamond patterns, as well as a reference monolithic non-upgradeable smart contract referred to as the Classic pattern.
For each pattern, the initial deployment of the notarization smart contract is referred to as Version 1. We will then proceed to two sequential upgrades of the application, a minor one (Version 2) and a major one (Version 3), to evaluate the gas costs. Features will be added as updates are made. Figure 2 demonstrates the upgrades for each pattern.
Figure 2: Evolution of the Notarization Application Across Versions by Patterns
Then the analysis turns to comparing the average gas cost of each function, for seeing which pattern is the less gas intensive when it comes to execution. Each function is accompanied by a test case that is run hundreds of times, to simulate real-world usage of this file notarization application over an extended period and to check the robustness of the code.
4.2 Implementation
The code will be written in Solidity, the main language used to develop decentralized application on EVM blockchains. For the experiment, the Universal Upgradeable Proxy Standard (UUPS) Proxy pattern is used because it is the recommended standard at the time of writing by OpenZeppelin [7]. For this study, the Diamond pattern implemented by SolidState is used. It utilizes Nick Mudge’s gas-efficient Diamond 2 model [29, 24].
4.2.1 Version 1: Basic Notarization Application
In the first version of the notarization use case, files are represented by a mapping between their name and their hash. The logic will contain the following functions.
  • addFile — A function that notarizes a file on the blockchain. It will modify the state of the contract.
  • getFileName — A view function that returns the name of the file.
  • getFileHash — A view function that returns the hash of the file.
  • compareHash — A pure function that compares two hashes passed as parameters. This operation is ideally performed off-chain, but it will be used here to demonstrate the gas cost of a pure function.
4.2.2 Version 2: Updatable File
In case of file modifications, a function for updating it on the blockchain is needed. The second version of this notarization application will contain the updateFile function, added on top of the previous version. This function updates the file hash of a notarized file and so modifies the state of the contract. This version can be seen as a minor update.
4.2.3 Version 3: Access Control
For security reasons, access control is mandatory, so that only the owner of a file can modify or delete it. This involves creating a File structure, containing the owner’s address, the hash of the file’s contents, the creation timestamp and the last modification timestamp. The previous mapping will be replaced by one between the file’s name and its File structure instance. Then, the addFile and updateFile functions are modified in order to work with this new File structure, and access control is added, where the new logic smart contract ensures that the caller interacts only with his own files. This version can be seen as a major update.
Unit tests are written using Foundry, a smart contract development framework built in Rust [19]. This toolkit is chosen because it compiles and executes faster than other well-known frameworks such as Truffle and Hardhat. Moreover, it eliminates the need for an additional programming language for writing tests and scripts, unlike Truffle and Hardhat with JavaScript, and generates easily gas reports per smart contract function.
The addFile function, in particular, has three test cases with different configurations for the file name parameter to ensure comprehensive coverage. As mentioned earlier, these test cases are conducted repeatedly, with each iteration in the three different configurations of the addFile file name parameter following these criteria: (i) The file name length is incremented by one character, (ii) the file name length remains constant, but the last character varies, (iii) both the file name and hash are unchanged. Foundry autonomously simulates an EVM blockchain environment during test execution. It deploys the contracts and carries out the testing scenarios within this emulated setting. The testing utilizes the default configuration of this local blockchain.
The entire source code, tests and results here are available in the accompanying source code repository. [^4] https://github.com/opixelum/proxy-diamond-patterns-gas-analysis
4.3 Results Retrieval
Through these unit tests, Foundry produces gas reports that provide insights into the gas consumption for each function, alongside the gas costs associated with deployment and the size of the contracts. The cost of function calls is quantified in gas units, and the contract size is measured in bytes.
In addition, the framework furnishes complete stack traces of the calls, detailing the gas consumption for each operation. These traces are utilized to construct charts that capture the gas cost of every individual function call. This level of detail supplements gas reports, which typically provide only a summary of costs, including the minimum, average, median, and maximum values.
5 Evaluation
This section presents the gas costs evaluation of the Proxy and Diamond smart contract patterns, against a baseline built using the classic pattern in the context of an app deployment and upgrade  [^5] All results can be found here: https://github.com/opixelum/proxy-diamond-patterns-gas-analysis/tree/main/data.
5.1 Gas Cost During Upgradeable Smart Contract Deployment and Upgrades
Figure 3: Deployment Cost by Pattern & Version.
We first aim at comparing the gas costs during the deployment (Version 1) and upgrades (Version 2 and 3) of the file notarization application implemented with the three different patterns. For the Classic and Proxy pattern, a new contract is deployed for each version. For the Proxy, the proxy pointer is updated in the smart contract implementation. For the Diamond pattern, in Version 2, only facet is changed and Version 3, all facets are changed. Figure 3 presents the deployment costs, in gas units, of each version by pattern. The cost of deployment increases with each version, except for the Diamond pattern, where the initial cost of deployment is significant. In total, the classic pattern emerges requires 1,614,545 gas units, while the Proxy and Diamond patterns consume approximately 2.6 times more, with 4,343,104 and 4,123,977 gas units respectively. The classic pattern appears the most gas efficient. The increased consumption at deployment of the Proxy and Diamond patterns relates to the need to deploy more contracts compared to the classic pattern: two for the Proxy pattern, and four for our example with the Diamond pattern.
5.2 Gas cost during upgradeable smart contract execution: study of all functions
Figure 4: Average Function Cost by Pattern for Version 2
The analysis now turns to comparing the average gas cost of each function, for seeing which pattern is the most gas efficient when it comes to execution. Figure 4 depicts the gas costs observed for Version 2 averaged per function. It is to note that we observe the same gas costs behavior in Version 1 and 3 (c.f., online gas reports). The addFile function is the costliest, followed by the updateFile function. The other functions require less gas. There is no significant difference in gas cost among the three patterns. The Proxy pattern is slightly more expensive than the classic pattern, and the Diamond pattern is slightly more expensive than the Proxy pattern. The most important difference is for compareHashes being 300 times cheaper than addFile for the classic pattern.
The difference in gas costs between compareHashes and addFile is due to the type of operations done in the functions. The reason is that the addFile does two writings in the storage, whereas the updateFile does only one. The other functions only do reads in the storage, which cost less gas than writes.
The classic pattern does not delegate any call to another contract, whereas the Proxy and Diamond patterns do, so it adds a cost. Then, when calling a diamond, it needs to do more operations than a proxy to find the right facet to call. This issue is not present with the Proxy pattern, as there is only one implementation contract. Now, the differences between the view and pure functions are examined. getFile... functions are view functions, which means that they only do a state lookup. This lookup can cost up to 2100 gas for a cold storage load (when reading a variable for the first time) [1]. compareHashes is a pure function, which means that it does not do any state lookup or modification. It only does some computations on given parameters, blockchain’s global variables or not. This is why it requires less gas than getFile... functions in the classic pattern.
5.3 Gas cost during upgradeable smart contract execution: more in-depth study of the addFile function
Figure 5: Gas Cost of the addFile Function by File Name Configuration
To gain a deeper understanding of the gas cost implications, the addFile function is subjected to three specific test cases, each with different configurations for the file name parameter to provide a more detailed analysis of how various input scenarios impact gas consumption. Figure 5 provides the results of the experiment. For the classic pattern, gas cost increases after each 32 characters, and remains the same for each iteration. Gas cost is expensive for the first iteration, then has an important drop for all the other iterations (it stays the same for each remaining iteration). Results are similar for the Proxy and Diamond patterns. The results for the first configuration can be explained bt the storage slot that is 32 bytes long. When a storage slot is full, the EVM needs to use another storage slot, which has a cost. That is why in the second configuration, the gas cost is the same for each iteration, since the length of the file name is the same. But because the last character is different, the EVM needs to write in the storage, which has a cost. Finally, the last configuration is way cheaper than the previous one after the first iteration because the EVM does not need to edit the storage, since the file name is the same.
5.4 Summary of Comparative Test Findings
In reviewing the data obtained from the comparative analyses, it is evident that the deployment phase exhibits the most significant variation among the different patterns, primarily due to the fluctuating number of smart contracts deployed. During execution, the gas costs are relatively similar across the patterns, though there is a marginal escalation in costs when transitioning from the classic architecture to the Proxy pattern, and subsequently from the Proxy to the Diamond pattern. However, this incremental increase in gas costs may be considered negligible, depending on the specific requirements of the software.
6 Theory Generation
6.1 Proxy and Diamond Patterns
We propose to summarize and consolidate the Proxy and Diamond blockchain patterns by following the Alexandrian form format proposed by Christopher Alexander [3].
6.1.1 Proxy Pattern
  • Summary - The Proxy Pattern facilitates upgradability in smart contracts through a Proxy contract pointing to the latest version of a logic contract.
  • Context - A smart contract must be upgraded due to evolving requirements and potential improvements [9].
  • Problem - Traditional smart contracts lack the ability to be updated without manual storage migration, posing challenges in addressing vulnerabilities, enhancing functionality, and adapting to changing circumstances.
  • Forces
    • Immutability vs. Upgradability - Smart contracts on blockchain platforms are traditionally immutable once deployed, ensuring security but hindering adaptability.
    • Gas Costs vs. Flexibility - The balance between minimizing gas costs and providing flexibility for contract upgrades is a critical challenge.
    • Trust vs. Transparency - Establishing trust in the upgrade process while maintaining transparency is a delicate balance.
  • Solution - The Smart Contract Proxy Pattern introduces a proxy contract as an intermediary layer. This proxy delegates calls to a logic contract, allowing seamless upgrades by deploying new logic contracts and updating the proxy to point to the latest version, hence allowing for dynamic updates.
  • Consequences
    • Benefits:
        Upgradability: Enables systematic and non-disruptive upgrades by employing a distinct proxy contract that separates logic and state. Manual data migration is not required, as the storage is maintained within the proxy contract. This allows the logic contract to be updated without altering the state-holding proxy contract.
        Seamless Interaction with Existing State: Maintains the contract’s address and state during upgrades when the logic contract in the Proxy pattern is upgraded. The new version can seamlessly interact with the existing state, minimizing the need for explicit data migration.
        Ease of implementation: Several libraries offer an easy integration of this pattern into smart contracts, encapsulating the proxy management (e.g., Open Zepplin and EIP-897 implementations)
    • Drawbacks:
        Proxy Management Functions Complexity: Requires dedicated proxy management functions, contingent on the type of proxy used (transparent or UUPS). This introduces an extra layer of complexity in the development process, demanding specific knowledge and consideration during implementation.
        Compatibility Maintenance: Ensuring compatibility between different contract versions in the Proxy pattern requires meticulous consideration. Preserving consistent function selectors and storage layouts becomes crucial for a smooth transition during upgrades.
        Limited Direct Function Visibility: Indirect access to the logic functions visibility. Users only have a direct access to the proxy administration functions, necessitating consultation of accompanying documentation to identify callable functions. This may potentially introduce usability challenges for end-users.
  • Related Patterns - Diamond Pattern and upgradeable pattern variations [8, 4]
  • Known Uses
    • EIP-897: ERC-Proxy Standard: An Ethereum Improvement Proposal outlining a standardized implementation of proxy contracts[23].
    • OpenZeppelin’s Proxy Patterns: Practical implementations and documentation in popular smart contract development frameworks[4].
    • Compound and USDC: Respectively a DeFi protocol and a stable coin leveraging the Proxy Pattern for upgrades
6.1.2 Diamond Pattern
  • Summary - The Diamond Pattern addresses the challenge of contract size limitations and enhances maintainability and versioning by employing multiple implementation contracts.
  • Context - Need for a solution to improve maintainability in large contracts.
  • Problem - Traditional approaches face challenges in managing contract size and versioning effectively, limiting the scalability and maintenance of smart contracts. The need for more scalable and maintainable smart contract structures is evident.
  • Forces
    • Contract Size vs. Modularity - Balancing the need for compact contract sizes with the demand for modular, well-organized code structures poses a challenge.
    • Maintainability vs. Simplicity - Achieving improved maintainability without introducing unnecessary complexity is an ongoing struggle.
    • Scalability vs. Consistency - Scaling smart contract applications can hinder consistency as it requires accommodating versioning and updates.
  • Solution - The Diamond Pattern introduces a structure where a proxy can point to multiple logic contracts. The Diamond pattern deployment involves the deployment of all contracts (diamond and facets), retrieving facets function selectors, and leveraging both to implement a diamond cut.
  • Consequences
    • Benefits:
        Structured and Modular design: Allows for multiple logic contracts (facets) which facilitate large contracts code organization.
        Flexible maintenance: Easy addition, replacement, or removal of facets, facilitating smoother upgrades and interactions with existing data, as well as scalable versioning.
        Compilation time for facet updates: The modular facet-based code organization shortens compilation time, as editing a facet only requires the compilation of that specific facet.
    • Drawbacks:
        Implementation Complexity: Complexity in smart contract variables storage management and facet-library interactions. This increased complexity demands specific knowledge and careful consideration during the implementation phase as there is currently a lack of supporting implementation libraries.
        Complexity in managing multiple logic contracts: The Diamond pattern’s enhanced flexibility comes at the cost of increased complexity, necessitating meticulous planning and testing to ensure smooth data interaction and integrity across different contract versions. The intricacies of managing multiple logic contracts require careful consideration to avoid unintended consequences during upgrades.
  • Related Patterns - Proxy Pattern and upgradeable pattern variations [8, 4]
  • Known Uses
    • Ethereum Improvement Proposal EIP-2535 [8]
    • Aavegotchi: a Non-Fungible Token (NFT) based gaming protocol
    • GeoWeb: a dApp managing digital land property rights using NFTs
6.2 Choosing between Proxy and Diamond Patterns
Figure 6: Decision model for Upgradeable Smart Contract Pattern Usage
Figure 6 proposes a decision model to orient between a Proxy or a Diamond pattern in the design stage of an upgradeable smart contract. It follows the design goal decision model introduced by Xu  [32] where decisions are modeled using a logical BPMN flow. More precisely, arrows, logical gateways, and functional and non-functional properties orient the decision path.
The choice of smart contract pattern largely depends on the need for upgradability. For scenarios without upgradability needs, the classic pattern is preferable due to its straightforward development process, relying on contract inheritance and library imports. However, upgrades in the classic pattern often involve complex and resource-intensive data migration. This pattern also poses challenges in communicating new contract addresses to users, potentially affecting user experience. For extensive upgradeable features, the Diamond Pattern is recommended. Its modular nature allows for easy addition or removal of facets, reducing compilation time. However, it is less cost-effective initially compared to the Proxy pattern and requires in-depth knowledge of smart contract storage and facet-library management. The Proxy pattern is advised for limited code sizes or infrequent upgrades. It simplifies development and integrates easily with libraries like OpenZeppelin’s. This pattern enhances upgradability by separating logic and state, reducing the need for data migration. But, it offers less flexibility and modularity compared to the Diamond Pattern and demands careful consideration to maintain compatibility across versions.
In conclusion, while the classic pattern excels in execution ease, the Proxy and Diamond patterns provide a more manageable framework for upgrades, simplifying contract interactions for users during updates.
7 Discussion and Conclusion
In conclusion, this comprehensive study delves into the intricacies of upgradeable smart contract design patterns—a critical facet of contemporary blockchain applications. The comparative analysis specifically focuses on the gas costs associated with deploying, using, and upgrading decentralized applications using two prominent upgradeable patterns: the Proxy and Diamond patterns.
Each pattern unfolds with distinct strengths and weaknesses, delineating its applicability across diverse scenarios. The Diamond pattern exhibits modularity countered by higher initial costs, and the Proxy pattern strikes a balance between the classical approach (without proxy) and the Diamond one.
To generalize these initial findings, the study advocates for extending experiments beyond the notarization scenario used as a comparison baseline for this paper. Essential to this endeavor is the necessity for additional experiments encompassing diverse blockchain networks and pattern libraries to extrapolate and validate the findings. Moreover, a replication of the study on alternative upgradeable patterns would also provide a more comprehensive understanding of upgradeable smart contract patterns.
Despite these considerations, real-world implementations in projects like Compound, USDC, GeoWeb, and Aavegotchi underscore the substantial impact of the Proxy and Diamond patterns in the blockchain domain. These patterns serve as facilitators of flexible, upgradeable, and scalable smart contracts, showcasing their adaptability across a spectrum of blockchain applications. In the dynamic landscape of the blockchain industry, the strategic selection and adept implementation of these design patterns emerge as pivotal determinants of success and innovation in decentralized applications.
Looking ahead, future research should delve into the dynamic nature of these patterns, potentially exploring hybrid upgradability models or introducing novel paradigms in smart contract development. Additionally, a deeper exploration into blockchain sharding and layer 2 solutions holds promise for offering valuable insights into the scalability of these patterns in diverse blockchain ecosystems.